Mr. King purchased 35 new cars at $15,000 each. He sold 20% of them for $19,000 each and the remainder for $21,000. What was his profit
Answer by milo1954(10)

total purchase 35*15,000=525,000
sale
20% of 35 =7 sold for 19,000
7*19,000=133,000
remainder
80% of 35 = 28 sold for 21,000
28*21,000=588,000
total sold=133,000+588,000
total sold=721,000
then
profit=total sold - total purchase
profit=721,000-525,000
profit=196,000